CH3NH2 + H2O <-> CH3NH3+ + OH-

Ka = 2.33x10^-11

Given the CH3NH2 concentration is 1.5M, what is the equilibrium concentration of Ch3NH2, CH3NH3+, OH-?


1
Expert's answer
2021-07-19T04:05:07-0400

2.33× 10-11 = 1.5 × x

= 2.33/1.5

× = 1.533M